Project

General

Profile

June 2014 Minutes

Attending: Susan, Nick, Dom, Michael, Fernanda, Mat, Gavin, Pavan, Nitin, Siva:

06/25/2014 Minutes
h2. 06/09/2014 Minutes

Detector Performance:

Lost 8 hrs due sync - timing system issue. (Experts fixing)

Jump in the hits/spill from 2pm 6/17-7 am 6/18: Associated with numi trigger rate? DCM on/off in hall?

PE distribution - The foot at the high end of the distribution in almost gone in the week plot... as compared with the day of month version. Changing thresholds on the detector?

Neutrino scanning- Some interesting events observed. Michael will report on Friday.

-----------------
NDOS, ND - Off

-----------------
Watchdog page - Have been rock solid recently. No edge gap any more.

Good data selection:
Joao is ready to start copying his good data plots. Nitin will need publish this.

Siva- Nearline reprocessing
- (Michael) We will do it eventually.
- Work to run DAQ to Raw Nearline, then Nearline regular in single job
- Email nova data quality

Pavan - Asked about cosmic trigger rates double line. It's related to discrete nature of data.